# Parameter inference of microlensed gravitational waves using neural spline flows

### v1

@@ IV Conclusion and outlook
In this work, we investigated the applicability of the NSF network in parameter inference for microlensed GW events. Specifically, we trained an NSF network to infer the 13-dimensional parameters describing the lens and source in the point-mass lens model. The network infers parameters $M_{\rm L}$, $D_{\rm L}$, $Z_{\rm L}$, $\mathcal{M}_{\rm c}$, $D_{\rm L}$, $Z_{\rm S}$, and $\iota$ not only much faster than Bilby dynesty but also with slightly better accuracy. Although the accuracy of inferring parameters $q$, $dec$, $ra$, $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$ is not ideal, it can still rapidly narrow down the estimated range of these parameters. Moreover, the NSF network demonstrates good generalization for spin parameters not included in the training, indicating that it may be capable of adapting to a wider variety of events.
In conclusion, this study demonstrates that NSFs can efficiently complete the parameter inference task for high-dimensional parameter lensed GW events. Compared with the traditional Bayesian method, NSFs shows significant advantages in computational efficiency, with its average running speed being four orders of magnitude faster than that of Bilby dynesty. This characteristic makes it promising to provide strong support for the low-latency search of lensed GW events.
@@
In the future, we will further investigate why the NSF network is not accurate enough in inferring parameters such as $q$, $dec$, $ra$, $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$. This may involve improving the architecture of the NSF network itself. We will also explore the use of the NSF network’s ability to quickly narrow down the inference range of these parameters, and use the network’s posterior estimates of these parameters as the prior for bilby to complete the entire parameter inference process Williams et al. (2021); Nerin et al. (2025). This will help us explore the combined use of the NSF network and traditional parameter inference methods. Additionally, for most lensed GW signals, both strong lensing and microlensing effects are present simultaneously. Some investigations have shown that strong lenses make microlensing effects easier to detect Diego (2020); Mishra et al. (2021); Cheung et al. (2021); Seo et al. (2022), and further research on the parameter inference effect under this combined effect will be more practical and valuable.

### v2

@@ IV Conclusion and outlook
In this work, we investigated the applicability of the NSF network in parameter inference for microlensed GW events. Specifically, we trained an NSF network to infer the 11-dimensional parameters describing the lens and source in the point-mass lens model. For a randomly selected event, the network infers parameters $M_{\rm L}$, $D_{\rm L}$, $y$, $\mathcal{M}_{\rm c}$, $D_{\rm L}$, $q$, $\delta$, $\alpha$ and $\iota$ not only much faster than Bilby dynesty, but also did not differ too much in accuracy. Although the accuracy of inferring parameters $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$ is not ideal, it can still rapidly narrow down the estimated range of these parameters. Moreover, the NSF network demonstrates good generalization for spin parameters not included in the training, indicating that it may be capable of adapting to a wider variety of events.
In conclusion, this study demonstrates that NSFs can significantly enhance the search efficiency within the parameter space and can efficiently complete the parameter inference for microlensed GW events.Compared with the traditional Bayesian method, NSFs’ average inference speed in a 12-dimensional parameter space is four orders of magnitude faster than that of Bilby dynesty. This makes it promising to provide strong support for low-latency searches of lensed GW events.
@@
In the future, we will further investigate why the NSF network is not accurate enough in inferring parameters such as $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$. This may involve improving the architecture of the NSF network itself. We will also explore the use of the NSF network’s ability to quickly narrow down the inference range of these parameters, and use the network’s posterior estimates of these parameters as the prior for bilby to complete the entire parameter inference process Williams et al. (2021); Nerin et al. (2025). This will help us explore the combined use of the NSF network and traditional parameter inference methods. Additionally, for most lensed GW signals, both strong lensing and microlensing effects are present simultaneously. Some investigations have shown that strong lenses make microlensing effects easier to detect Diego (2020); Mishra et al. (2021); Cheung et al. (2021); Seo et al. (2022), and further research on the parameter inference effect under this combined effect will be more practical and valuable.
DATA AVAILABILITY
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing,and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
SOFTWARE AVAILABILITY
The code used for simulating GW and data processing techniques is used by Pycbc. The code for lensing the simulated GW can be requested from the corresponding authors upon reasonable request.